Given the two factor product in Q= 150L^0.5 K^ 0.5, wage rate of labour = 50 and rental cost of capital = 40. Determine the amount of labour and capital that will minimize the cost of producing 1118 units of output?

1
Expert's answer
2022-02-15T11:18:07-0500

"Q=150L^{0.5}K^{0.5}"

"C=50L+40K"

Fitting in the langragian equation:

"l=50L+40K-h(1118-150L^{0.5}K^{0.5})"

Differentiate the langragian equation with respect to L and K:

"\frac{dl}{dL}=50-h75^{-0.5}K^{0.5}=0......(i)"

"\frac{dl}{dK}=50-h75L^{0.5}K^{-0.5}=0.....(ii)"

Equating equation (i) and (ii):

"50-h75L^{-0.5}K^{0.5}=50-h75^{0.5}K^{-0.5}"

Like terms on both sides cancel each other:

"L^{-0.5}K^{0.5}=L^{0.5}K^{-0.5}"

"\frac{K^{0.5}}{L^{0.5}}=\frac{L^{0.5}}{K^{0.5}}"

Therefore "K=L"

Replacing K and Q in

"Q=150L^{0.5}K^{0.5}"

We get:

"1119=150L^{0.5}L^{0.5}"

"1118=150L^{2}"

"L=2.7"

Replacing L we get:

"1118=150\u00d7(2.7)^{0.5}K^{0.5}"

"4.54=\u221aK"

"K=2.13"